What is the center of the hyperbola whose equation is [(x+2)2 / 9] - [(y-2)2 / 16] = 1?
Answer by lwsshak3(11628) 
 What is the center of the hyperbola whose equation is [(x+2)2 / 9] - [(y-2)2 / 16] = 1?
**
(x+2)^2/9-(y-2)^2/16 = 1
This is an equation of a hyperbola with horizontal transverse axis of the standard form:
(x-h)^2/a^2-(y-k)^2/b^2=1, with (h,k) being the (x,y) coordinates of the center.
For given equation:
center: (-2,2)
|
|
|
| |